OpenStudy (anonymous):

Post here all your algebra problems, but no cos sin or tan. I'll try to solve it for you.

OpenStudy (anonymous):

\[5^{15}\times \left[\begin{matrix}2 & 5 \\ 12 & 9\end{matrix}\right]\]

OpenStudy (anonymous):

5^15 x [(2*9)-(12*5)]? Linear Algebra is not my strongest atribute :P

OpenStudy (loser66):

How about modern algebra? I need a lot of help:)

OpenStudy (anonymous):

Give me the questions I'll try to do it i'm bored :)

OpenStudy (loser66):

Let I, J be 2 ideals in a ring R. Show that \[IJ=\{a_1b_1+a_2b_2+....+a_sb_s~~| ~~s\geq 1, a_i\in I~~ and~~ b_i\in J\}\] is an ideal in R and \(IJ\subset I\cap J\)

OpenStudy (anonymous):

@Loser66 show offff :D :D

OpenStudy (anonymous):

Ow thats out of my knowledge, I thought algebra was solving for x in complex formulas xD, we call it different over here thought it was called algebra :P

OpenStudy (loser66):

@magepker728 you mean I show offff?? nope, I desperately need help

OpenStudy (anonymous):

What do you study?

OpenStudy (loser66):

Modern algebra

OpenStudy (anonymous):

Nice :)

OpenStudy (anonymous):

hell no bruh, uve help me before so i know a bit about ur knowledge lol

OpenStudy (anonymous):

I'm only an economic student, but I love solving for x in complex formulas xD

OpenStudy (anonymous):

ok here is an easy one..